1) At an air show, a jet plane has velocity components vx = 675 km/h and vy = 515 km/h at time 3.90 s and vx = 878 km/hand vy = 385 km/h at time 7.27 s

Part A) For this time interval, find the x component of the plane's average acceleration.

Part B) For this time interval, find the y component of the plane's average acceleration.

Part C) For this time interval, find the magnitude of its average acceleration.

Part D) For this time interval, find the direction of its average acceleration.
Answer #1

Given vx = 675 km/h and vy = 515 km/h

time t = 3.90 s = 3.9/(60x60)hrs

vx ' = 878 km/h

vy ' = 385 km/h

time t ' =7.27 s =7.27/(60x60) hrs

(A)..the x component of the plane's average acceleration ax = (vx ' - vx) /(t ' - t )

                         ax = (878-675)/[(7.27-3.9)/(60x60)]

                               = 216854.6 km/h 2

B) the y component of the plane's average acceleration ay = (vy '-vy)/(t'-t)

ay = (385-515)/[(7.27-3.9)/(60x60)]

       = -138872.4 km/h 2

C) the magnitude of its average acceleration a = \sqrt{}[ax 2+ay 2]

               a = 257510.1 km/h 2

D) Let the accleration makes an angle \theta with X axis then ,

          tan \theta = ay/ax

                  = -138872.4 / 216854.6

                  = -0.64

                \theta = -32.63 o